OpenStudy (clairebracken1234):

Given that $p\ge 7$ is a prime number, evaluate $$1^{-1} \cdot 2^{-1} + 2^{-1} \cdot 3^{-1} + 3^{-1} \cdot 4^{-1} + \cdots + (p-2)^{-1} \cdot (p-1)^{-1} \pmod{p}.$$

OpenStudy (clairebracken1234):

p >= 7*

jimthompson5910 (jim_thompson5910):

Let's try different values of p to see how the sum is expanded out I'm going to try p = 7 first since it's the smallest prime p allowed here \[1^{-1} \cdot 2^{-1} + 2^{-1} \cdot 3^{-1} + 3^{-1} \cdot 4^{-1} + \cdots + (p-2)^{-1} \cdot (p-1)^{-1} \pmod{p}.\] \[1^{-1} \cdot 2^{-1} + 2^{-1} \cdot 3^{-1} + 3^{-1} \cdot 4^{-1} + \cdots + (7-2)^{-1} \cdot (7-1)^{-1} \pmod{7}.\] \[1^{-1} \cdot 2^{-1} + 2^{-1} \cdot 3^{-1} + 3^{-1} \cdot 4^{-1} + \cdots + 5^{-1} \cdot 6^{-1} \pmod{7}.\] \[1^{-1} \cdot 2^{-1} + 2^{-1} \cdot 3^{-1} + 3^{-1} \cdot 4^{-1}+ 4^{-1} \cdot 5^{-1} + 5^{-1} \cdot 6^{-1} \pmod{7}.\] \[1 \cdot 4 + 4 \cdot 5 + 5 \cdot 2+ 2 \cdot 3 + 3 \cdot 6 \pmod{7}.\] \[4+20+10+6+18 \pmod{7}.\] \[58 \pmod{7}.\] \[2 \pmod{7}.\]

jimthompson5910 (jim_thompson5910):

That's a specific case (p = 7). I'm still trying to think of how to do this in general for any prime p >= 7

jimthompson5910 (jim_thompson5910):

Did your professor give you any hints or formulas to use? I'm still stuck. I have a feeling this uses fermat's little theorem somehow. I'm not sure though.

OpenStudy (clairebracken1234):

Nope haha its an Alcumus problem on AOPS It's supposed to use fermat's little theorem but im sorta confused @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Let me try p = 7 a different way \[1^{-1} \cdot 2^{-1} + 2^{-1} \cdot 3^{-1} + 3^{-1} \cdot 4^{-1} + \cdots + (p-2)^{-1} \cdot (p-1)^{-1} \pmod{p}.\] \[1^{-1} \cdot 2^{-1} + 2^{-1} \cdot 3^{-1} + 3^{-1} \cdot 4^{-1} + \cdots + (7-2)^{-1} \cdot (7-1)^{-1} \pmod{7}.\] \[1^{-1} \cdot 2^{-1} + 2^{-1} \cdot 3^{-1} + 3^{-1} \cdot 4^{-1} + \cdots + 5^{-1} \cdot 6^{-1} \pmod{7}.\] \[1^{-1} \cdot 2^{-1} + 2^{-1} \cdot 3^{-1} + 3^{-1} \cdot 4^{-1}+ 4^{-1} \cdot 5^{-1} + 5^{-1} \cdot 6^{-1} \pmod{7}.\] \[(1*2)^{-1} + (2*3)^{-1} + (3*4)^{-1}+ (4*5)^{-1} + (5*6)^{-1} \pmod{7}.\] \[2^{-1} + 6^{-1} + 12^{-1}+ 20^{-1} + 30^{-1} \pmod{7}.\] \[2^{-1} + 6^{-1} + 5^{-1}+ 6^{-1} + 2^{-1} \pmod{7}.\] So that's a bit interesting. We have this symmetry going on. Let me try p = 11 next

jimthompson5910 (jim_thompson5910):

I'm going to skip a few steps to save time/space \[(1*2)^{-1} + (2*3)^{-1} + (3*4)^{-1}+ (4*5)^{-1} + (5*6)^{-1} + (6*7)^{-1}\\+ (7*8)^{-1} + (8*9)^{-1} + (9*10)^{-1} \pmod{11}.\] \[2^{-1} + 6^{-1} + 12^{-1}+ 20^{-1} + 30^{-1} + 42^{-1} + 56^{-1} + 72^{-1} + 90^{-1} \pmod{11}.\] \[2^{-1} + 6^{-1} + 12^{-1}+ 9^{-1} + 8^{-1} + 8^{-1} + 1^{-1} + 6^{-1} + 2^{-1} \pmod{11}.\] This has a bit of symmetry but not as much as the last on did. The 9 and 1 don't match up. I'll try p = 13 next

jimthompson5910 (jim_thompson5910):

p = 13 \[(1*2)^{-1} + (2*3)^{-1} + (3*4)^{-1}+ (4*5)^{-1} + (5*6)^{-1} + (6*7)^{-1} + (7*8)^{-1}\\+(8*9)^{-1} + (9*10)^{-1}+(10*11)^{-1} + (11*12)^{-1} \pmod{13}.\] \[2^{-1} + 6^{-1} + 12^{-1}+ 20^{-1} + 30^{-1} + 42^{-1} + 56^{-1} + 72^{-1} + 90^{-1}+110^{-1}+132^{-1} \pmod{13}.\] \[2^{-1} + 6^{-1} + 12^{-1}+ 7^{-1} + 4^{-1} + 3^{-1} + 4^{-1} + 7^{-1} + 12^{-1}+6^{-1}+2^{-1} \pmod{13}.\] Now this has perfect symmetry. I may have made a mistake with p = 11 somewhere?

OpenStudy (clairebracken1234):

so p = 13 works? @jim_thompson5910

OpenStudy (clairebracken1234):

i cant try entering 13 as the answer in alcumus and see the result and get the explanation @jim_thompson5910

jimthompson5910 (jim_thompson5910):

p = 13 has the same type of symmetry p = 7 does

jimthompson5910 (jim_thompson5910):

I'm still stumped on what the answer would be in general for any p I'm curious what explanation they give

OpenStudy (clairebracken1234):

i can try entering 13 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I doubt the answer is 13. That seems too easy, but yeah sure go ahead. Please post the explanation they give. I'm curious how to solve this one

OpenStudy (clairebracken1234):

haha it's incorrect..i have 2 chances so should i type in something random just to find the answer? my score will go down but its okay lol @jim_thompson5910 this is a really hard problem

jimthompson5910 (jim_thompson5910):

yeah I've never seen anything this tough before. I'm probably overthinking

OpenStudy (clairebracken1234):

As $p$ is a prime number, it follows that the modular inverses of $1,2, \ldots, p-1$ all exist. We claim that $n^{-1} \cdot (n+1)^{-1} \equiv n^{-1} - (n+1)^{-1} \pmod{p}$ for $n \in \{1,2, \ldots, p-2\}$, in analogue with the formula $\frac{1}{n(n+1)} = \frac{1}{n} - \frac{1}{n+1}$. Indeed, multiplying both sides of the congruence by $n(n+1)$, we find that $$1 \equiv n(n+1) \cdot (n^{-1} - (n+1)^{-1}) \equiv (n+1) - n \equiv 1 \pmod{p},$$as desired. Thus, \begin{align*}&1^{-1} \cdot 2^{-1} + 2^{-1} + 3^{-1} + 3^{-1} \cdot 4^{-1} + \cdots + (p-2)^{-1} \cdot (p-1)^{-1} \\ &\equiv 1^{-1} - 2^{-1} + 2^{-1} - 3^{-1} + \cdots - (p-1)^{-1} \pmod{p}.\end{align*}This is a telescoping series, which sums to $1^{-1} - (p-1)^{-1} \equiv 1 - (-1)^{-1} \equiv \boxed{2} \pmod{p}$, since the modular inverse of $-1$ is itself. @jim_thompson5910

OpenStudy (clairebracken1234):

answer is 2 somehow it didn't quite show up correctly on here i apologize @jim_thompson5910

jimthompson5910 (jim_thompson5910):

You have to change the dollar signs to \ ( and \ ) As \(p\) is a prime number, it follows that the modular inverses of \(1,2, \ldots, p-1\) all exist. We claim that \(n^{-1} \cdot (n+1)^{-1} \equiv n^{-1} - (n+1)^{-1} \pmod{p}\) for \(n \in \{1,2, \ldots, p-2\}\), in analogue with the formula \(\frac{1}{n(n+1)} = \frac{1}{n} - \frac{1}{n+1}\). Indeed, multiplying both sides of the congruence by \(n(n+1)\), we find that 1=n(n+1)·(n-1-(n+1)-1)=(n+1)-n=1(modp), as desired. Thus, \[\begin{align*}&1^{-1} \cdot 2^{-1} + 2^{-1} + 3^{-1} + 3^{-1} \cdot 4^{-1} + \cdots + (p-2)^{-1} \cdot (p-1)^{-1} \\ &\equiv 1^{-1} - 2^{-1} + 2^{-1} - 3^{-1} + \cdots - (p-1)^{-1} \pmod{p}.\end{align*}\] This is a telescoping series, which sums to \(1^{-1} - (p-1)^{-1} \equiv 1 - (-1)^{-1} \equiv \boxed{2} \pmod{p}\), since the modular inverse of \(-1\) is itself.

jimthompson5910 (jim_thompson5910):

Definitely a tough problem. I don't know if I would have come up with that.

OpenStudy (clairebracken1234):

Same haha it didn't drop my score that much tho bc it was such a hard one @jim_thompson5910

OpenStudy (kainui):

I think this is a recognizeable telescoping series and in mod p ensures that all the inverses exist I guess.

jimthompson5910 (jim_thompson5910):

I think the key lies with this part \[n^{-1} \cdot (n+1)^{-1} \equiv n^{-1} - (n+1)^{-1} \pmod{p}\] if you can see how that works, then the telescoping part is a bit more clear

OpenStudy (kainui):

Here's one way to the solution: \[\sum_{n=1}^{p-2} \frac{1}{n(n+1)} = \sum_{n=1}^{p-2} \frac{1}{n} -\frac{1}{n+1}= 1 - \frac{1}{p-1}\] Funny thing about \(p-1\) is that it is it's own inverse since \((-1)^2 =1\) \[p-1 \equiv -1\]\[(p-1)(p-1) \equiv 1\] So you can replace in the above, \[\frac{1}{p-1} = p-1\] and we have: \[1- (p-1) \equiv 2\]

OpenStudy (clairebracken1234):

thanks for your help nevertheless :) @jim_thompson5910 @Kainui